What is the shortest wavelength present in the Brackett series of spectral lines?


1
Expert's answer
2021-12-06T09:39:49-0500

Rydberg's formula for the Bracket series looks like this:


"\frac{1}{\lambda}=R(\frac{1}{4^2}-\frac{1}{m^2})" , where "R=1.0974\cdot10^{7}\ (m^{-1})" .


If "m\to\infin" then


"\frac{1}{\lambda}=1.0974\cdot10^{7}\cdot(\frac{1}{4^2}-\frac{1}{\infin})\to\lambda=1458\ (nm)" . Answer